# Jamming is a first-order transition with quenched disorder in amorphous materials sheared by cyclic quasistatic deformations

### v2

@@ Introduction.
The result of $\mu$ shall be interpreted with care. In general, (i) $\mu=0$ corresponds to uniformity of $Z_{i}$, (ii) $\mu>0$ to hyperuniformity with a vanishing $\chi_{\rm J}$ in the thermodynamic limit, and (iii) $\mu<0$ to hyperfluctuations with a diverging $\chi_{\rm J}$ in the thermodynamic limit that typically appears at the critical point in a second-order phase transition. However, the negative $\mu$ in Fig. 6f is not due to the criticality of the jamming transition. Here it is essential to consider the volume fluctuations. The negative $\mu$ is obtained from a $\varphi$-controlled setup in our CAQS simulations. If we instead select configurations around a constant pressure $P$, then the fluctuations become significantly smaller, and $\mu\approx 0$ (Fig. 6d). This observation is consistent with a previous study Ikeda (2023), suggesting that the large fluctuation of $Z$ near $\varphi_{\rm J}^{\rm\infty}$ in the $\varphi$-controlled protocol might be originated from the fluctuation of $\varphi_{\rm J}^{N}$. In fact, careful measurement gives $\mu\simeq 1$ in a $P$-controlled protocol at jamming, confirming hyperuniformity Hexner et al. (2018, 2019). Ref. Hexner et al. (2019) also compares an ensemble of subsystems cut out from large packings, to an ensemble of whole systems with the same volume under periodic boundary conditions (similar to the $P$-controlled ensemble considered in this study): the contact fluctuations in the former are larger than those in the latter, and the convergence of these two are expected to occur in extremely large systems that are not attainable in current simulations – Due to this pre-asymptotic effect, the hyperuniform exponent $\mu\simeq 1$ is only observable in the first ensemble, while in the second ensemble, the apparent exponent is close to $\mu\approx 0$, consistent with our $P$-controlled data in Fig. 6d. In short, although the hyperuniform exponent $\mu=1$ is not directly observed, our results do not contradict the recently reported hyperuniform behavior of contact distributions at jamming.
We emphasize that the finite-size effects of the jammed peak $p_{\rm J}(Z)$ contribute negligibly to the scaling of $F(\varphi,N)$, $\chi_{\rm dis}(\varphi,N)$ and $\chi_{\rm con}(\varphi,N)$. In other words, the first-order nature of the jamming transition is independent of how $Z_{i}$ is spatially distributed in jammed packings.
@@
In the scaling analysis presented above, $Z$ is treated as the essential order parameter. In principle, the same kind of scaling analysis can be applied to any physical quantity $A$ that varies discontinuously at jamming (e.g., $A$ can be the bulk modulus $B$ or the fraction of non-rattlers $f_{\rm NR}$). The distribution $p(A)$ should have the same form as Eq. (3), $p(A)=(1-F_{\rm J})\delta(A)+F_{\rm J}p_{\rm J}(A)$. With fragile states removed, generally $A=0$ below jamming because no contacts are formed (e.g., $B=0$ and $f_{\rm NR}=0$), and thus the unjammed peak is always a delta function. The fraction of jammed states $F_{\rm J}$ only depends on the constructed ensemble, which determines how the states are sampled, and thus its scaling form $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}[(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}]$ is independent of $A$. The jammed peak $p_{\rm J}(A)$ does depend on $A$, but as shown above, $p_{\rm J}(A)$ is irrelevant to the interested scaling behavior of $F_{\rm J},\chi_{\rm dis}$ and $\chi_{\rm con}$. Thus, our conclusion drawn from the scaling analysis should be robust and generic, independent of which parameter $A$ is chosen for the scaling analysis, as long as $A$ can reflect the discontinuous nature of jamming.
Conclusion.
@@
In previous rapid quench simulations, it is observed that the width $w$ of the jamming density distribution $P(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-\Omega}$ with $\Omega=0.55\pm 0.03$ in both 2D and 3D O’hern et al. (2003). This finite-size scaling depends on the total number of particles, $N$, rather than on the system length $\ell$, and the exponent $\Omega\approx 1/2$ is independent of the dimensionality within the numerical accuracy. An explanation, as suggested previously, is that jamming is a second-order transition with an upper critical dimension $d_{\rm u}=2$ Goodrich et al. (2012). Here we propose an alternative interpretation: jamming is a first-order transition with quenched disorder, and consequently the finite-size scaling is a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$, independent of the dimensionality.
It is of particular interest to reconcile the first-order transition established here under quasi-static shearing and the second-order transition observed in previous finite-rate rheology (see (iv) in Introduction). Conventionally, first- and second-order transitions coexist in gas-liquid systems, but not in liquid-crystalline solid systems. Because the liquid-crystalline solid transition is accompanied with spontaneous symmetry breaking, it cannot be a continuous type. However, there is no apparent spontaneous symmetry breaking during liquid-amorphous solid transitions (such as the jamming transition). Thus there is no fundamental reason to exclude the possibility of the coexistence of discontinuous and continuous transitions between a liquid and an amorphous solid. This study opens an avenue to unify first-order (this study) and second-order Olsson and Teitel (2007) jamming transitions. Another direction of the future work is to study the competition between the sample-to-sample fluctuations due to disorder and the fluctuations associated to the granular temperature Pöschel and Luding (2001); Luding et al. (2021), and examine the “thermal effects” on the nature of the first-order jamming transition. Similar studies on the thermal vestiges of the zero-temperature physics have been recently carried out in the driven RFIM at finite temperatures Yao and Jack (2023). It would be also interesting to examine if friction can alter the order of the jamming transition Song et al. (2008); Ramaswamy et al. (2023).
Methods.
@@
where $\epsilon=1$ is the energy unit, $r_{ij}$ the distance between particles $i$ and $j$, $\sigma_{ij}=\frac{\sigma_{i}+\sigma_{j}}{2}$ the mean diameter, and $\Theta(x)$ the Heaviside step function. We set unit particle mass, and the diameter of small particles as unit length.
@@
Acknowledgements.
We warmly thank Hajime Yoshino, Stephen Teitel, Bulbul Chakraborty, Olivier Dauchot, Gilles Tarjus for inspiring discussions. We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 11974361 and 12047503), from Chinese Academy of Sciences (Grants ZDBS-LY-7017 and KGFZD-145-22-13), and from Wenzhou Institute (Grant WIUCASQD2023009). In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.
@@ Acknowledgements.
Figure S3: Distribution of the coordination number in two dimensions for $\gamma_{\rm max}=1$. (a) Fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.8435$. (b) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid line represents the fitting according to Eq. (4),with two fitting parameters $u=0.051$ and $\sigma_{\varphi}=0.043$.
@@
Figure S5: Distribution of the coordination number in three dimensions. (a) Fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}=0.648$. (b) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm j}^{\infty})N^{1/2}$. The solid line represents the fitting according to Eq. (4), with two fitting parameters $u=0.077$ and $\sigma_{\varphi}=0.051$.

### v4

@@ Results
The result of $\mu$ shall be interpreted with care. In general, (i) $\mu=0$ corresponds to uniformity of $Z_{i}$, (ii) $\mu>0$ to hyperuniformity with a vanishing $\chi_{\rm J}$ in the thermodynamic limit, and (iii) $\mu<0$ to hyperfluctuations with a diverging $\chi_{\rm J}$ in the thermodynamic limit that typically appears at the critical point in a second-order phase transition. However, the negative $\mu$ in Fig. 6f is not due to the criticality of the jamming transition. Here it is essential to consider the volume fluctuations. The negative $\mu$ is obtained from a $\varphi$-controlled setup in our CAQS simulations. If we instead select configurations around a constant pressure $P$, then the fluctuations become significantly smaller, and $\mu\approx 0$ (Fig. 6d). This observation is consistent with a previous study Ikeda (2023), suggesting that the large fluctuation of $Z$ near $\varphi_{\rm J}^{\rm\infty}$ in the $\varphi$-controlled protocol might be originated from the fluctuation of $\varphi_{\rm J}^{N}$. In fact, careful measurement gives $\mu\simeq 1$ in a $P$-controlled protocol at jamming, confirming hyperuniformity Hexner et al. (2018, 2019). Ref. Hexner et al. (2019) also compares an ensemble of subsystems cut out from large packings, to an ensemble of whole systems with the same volume under periodic boundary conditions (similar to the $P$-controlled ensemble considered in this study): the contact fluctuations in the former are larger than those in the latter, and the convergence of these two are expected to occur in extremely large systems that are not attainable in current simulations – Due to this pre-asymptotic effect, the hyperuniform exponent $\mu\simeq 1$ is only observable in the first ensemble, while in the second ensemble, the apparent exponent is close to $\mu\approx 0$, consistent with our $P$-controlled data in Fig. 6d. In short, although the hyperuniform exponent $\mu=1$ is not directly observed, our results do not contradict the recently reported hyperuniform behavior of contact distributions at jamming. We emphasize that the finite-size effects of the jammed peak $p_{\rm J}(Z)$ contribute negligibly to the scaling of $F(\varphi,N)$, $\chi_{\rm dis}(\varphi,N)$ and $\chi_{\rm con}(\varphi,N)$. In other words, the first-order nature of the jamming transition is independent of how $Z_{i}$ is spatially distributed in jammed packings.
@@
In the scaling analysis presented above, $Z$ is treated as the essential order parameter. In principle, the same kind of scaling analysis can be applied to any physical quantity $A$ that varies discontinuously at jamming (e.g., $A$ can be the bulk modulus $B$ or the fraction of non-rattlers $f_{\rm NR}$). The distribution $p(A)$ should have the same form as Eq. (3), $p(A)=(1-F_{\rm J})\delta(A)+F_{\rm J}p_{\rm J}(A)$. With fragile states removed, generally $A=0$ below jamming because no contacts are formed (e.g., $B=0$ and $f_{\rm NR}=0$), and thus the unjammed peak is always a delta function. The fraction of jammed states $F_{\rm J}$ only depends on the constructed ensemble, which determines how the states are sampled, and thus its scaling form $F_{\rm J}(\varphi,N)=\mathcal{F}_{\rm J}[(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}]$ is independent of $A$. The jammed peak $p_{\rm J}(A)$ does depend on $A$, but as shown above, $p_{\rm J}(A)$ is irrelevant to the interested scaling behavior of $F_{\rm J},\chi_{\rm dis}$ and $\chi_{\rm con}$. Thus, our conclusion drawn from the scaling analysis should be robust and generic, independent of which parameter $A$ is chosen for the scaling analysis, as long as $A$ can reflect the discontinuous nature of jamming. If one instead chooses a continuous variable, such as the pressure $P$, then the above scaling analysis would not work.
Discussion
@@
In previous rapid quench simulations, it is observed that the width $w$ of the jamming density distribution $P(\varphi_{\rm J}^{N})$ scales as $w\sim N^{-\Omega}$ with $\Omega=0.55\pm 0.03$ in both 2D and 3D O’Hern et al. (2003). This finite-size scaling depends on the total number of particles, $N$, rather than on the system length $\ell$, and the exponent $\Omega\approx 1/2$ is independent of the dimensionality within the numerical accuracy. An explanation, as suggested previously, is that jamming is a second-order transition with an upper critical dimension $d_{\rm u}=2$ Goodrich et al. (2012). Here we propose an alternative interpretation: jamming is a first-order transition with quenched disorder, and consequently the finite-size scaling is a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$, independent of the dimensionality. We expect that the same finite-size scaling applies to the steady state configurations generated by uniform simple shear.
It is of particular interest to reconcile the first-order transition established here under quasi-static shearing and the second-order transition observed in previous finite-rate rheology (see (iv) in Introduction). Conventionally, first- and second-order transitions coexist in gas-liquid systems, but not in liquid-crystalline solid systems. Because the liquid-crystalline solid transition is accompanied with spontaneous symmetry breaking, it cannot be a continuous type. However, there is no apparent spontaneous symmetry breaking during liquid-amorphous solid transitions (such as the jamming transition). Thus there is no fundamental reason to exclude the possibility of the coexistence of discontinuous and continuous transitions between a liquid and an amorphous solid. This study opens an avenue to unify first-order (this study) and second-order Olsson and Teitel (2007) jamming transitions.
In this study, we ignore thermal fluctuations. An interesting direction of the future work is to study the competition between the sample-to-sample fluctuations due to disorder and the fluctuations associated to the granular temperature Pöschel and Luding (2001); Luding et al. (2021), and examine the “thermal effects” on the nature of the first-order jamming transition. Similar studies on the thermal vestiges of the zero-temperature physics have been recently carried out in the driven RFIM at finite temperatures Yao and Jack (2023). It would be also interesting to examine if friction can alter the order of the jamming transition Song et al. (2008); Ramaswamy et al. (2023).
Methods
@@
where $\epsilon=1$ is the energy unit, $r_{ij}$ the distance between particles $i$ and $j$, $\sigma_{ij}=\frac{\sigma_{i}+\sigma_{j}}{2}$ the mean diameter, and $\Theta(x)$ the Heaviside step function. We set unit particle mass, and the mean diameter as the unit length.
@@
Data availability
Source data are provided with this paper.
Code availability
The simulation program is available at https://github.com/dengpan-cn/Library-of-Athermal-Quasi-static-Shear-Simulation-of-Granular-Matter.
Acknowledgments
We warmly thank Hajime Yoshino, Stephen Teitel, Bulbul Chakraborty, Olivier Dauchot, Gilles Tarjus, Yujie Wang for inspiring discussions. We acknowledge financial support from NSFC (Grants 12161141007, 11935002, 11974361 and 12047503), from Chinese Academy of Sciences (Grants ZDBS-LY-7017 and KGFZD-145-22-13), and from Wenzhou Institute (Grant WIUCASQD2023009). In this work access was granted to the High-Performance Computing Cluster of Institute of Theoretical Physics - the Chinese Academy of Sciences.
Author Contributions Statement
All authors contributed equally to this work.
Competing Interests Statement
The authors declare no competing interests.
@@ Acknowledgments
Figure S3: Jamming fraction in two dimensions for $\gamma_{\rm max}=1$. (a) Fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}={0.8435(1)}$. (b) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm J}^{\infty})N^{1/2}$. The solid line represents the fitting according to Eq. (4),with two fitting parameters $u={0.051(1)}$ and $\sigma_{\varphi}={0.043(1)}$.
@@
Figure S5: Constant-pressure shear simulations in two dimensions. (a) Stress $\sigma$ as a function of the strain $\gamma$ in simple athermal quasistatic shear, under a constant pressure $P=10^{-5}$. The yield strain is $\gamma_{\rm Y}\approx 0.1$ (arrow), beyond which the curve becomes nearly a plateau. (b) $\varphi$ as a function of the number of cycles $t$ in cyclic athermal quasistatic shear (in one cycle $\gamma$ varies as $0\rightarrow\gamma_{\rm max}\rightarrow-\gamma_{\rm max}\rightarrow 0$), with a fixed $P=10^{-5}$. The data of $\gamma_{\rm max}=0.05$ are fitted to a logarithmic form (dashed line), $\varphi(t)=\varphi_{\rm f}-\frac{\varphi_{\rm f}-\varphi_{0}}{1+B\log(1+t/\tau)}$, as proposed in Ref. [74]. Here $\varphi_{0}=0.8405$ is the packing fraction of un-strained initial systems; $\varphi_{\rm f}=0.8516(5)$, $B=0.61(15)$ and $\tau=0.60(18)$ are fitting parameters. The error bar represents the standard error of the mean.
Figure S6: Jamming fraction in three dimensions. (a) Fraction of jammed states $F_{\rm J}$ as a function of $\varphi$ for a few different $N$. The intersection of curves gives $\varphi_{\rm J}^{\infty}={0.6479(2)}$. (b) The data points of $F_{\rm J}$ with different $N$ collapse as a function of $(\varphi-\varphi_{\rm j}^{\infty})N^{1/2}$. The solid line represents the fitting according to Eq. (4), with two fitting parameters $u={0.077(2)}$ and $\sigma_{\varphi}={0.051(1)}$.